# Current validation practice undermines surgical AI development

### v1

@@ 7. Conflicts of Interest
| P2.6 | Non-suitability of hyperparameters from other domains | Hyperparameters or default metric configurations adopted from other domains (e.g., general computer vision) may not align with the specific underlying clinical needs. For example, in object detection, a high Intersection over Union (IoU) threshold may be appropriate in general computer vision tasks, but in surgical applications, lower IoU thresholds might be more suitable when only rough localization of objects (e.g., polyps) is needed [tran2023sources]. Similarly, default hyperparameters for event detection in action recognition may not account for the variability in surgical workflows. In marker-less tool tracking (e.g., [hein2025next]), for example, thresholds that define acceptable accuracy in everyday scenarios (e.g., a few millimeters) may be insufficient in surgical contexts, where sub-millimeter precision can be clinically critical. | [tran2023sources, ramesh2023dissecting, hein2025next] | Extended Data Fig. 8b |
@@
| P2.7 | Intrinsic limitations of individual metrics | The used metrics harbor pitfalls related to their individual mathematical properties. Even if metrics are well-aligned with the clinical task, they may exhibit problematic behaviors such as sensitivity to class imbalance, non-linearity, or lack of interpretability. The suitability of any given metric should thus be analyzed in light of its known limitations. For example, accuracy may produce misleading values in highly imbalanced data sets. | [reinke2021common, reinke2024understanding] | |
@@
| P3.1 | Non-independence within the test set | Failure to account for non-independence within the test set can yield misleading conclusions. Non-independence can occur when multiple frames from the same surgical video are included, or when data from the same patient, procedure, or institution are used. Aggregation, statistical analysis and reporting need to account for this lack of independence rather than assuming independent samples. For example, if frames from the same patient are used for validation, performance metric values may appear artificially high due to strong temporal correlation. Other possible consequences include biased statistical estimates, such as underestimated uncertainty or distorted means. | [kostiuchik2024surgical, funke2023metrics, kirtac2022surgical] | Fig. 4 |
@@
| P3.2 | Clinically uninformative aggregation | Performance metrics are often simply aggregated, without accounting for the varying importance of different surgical phases or time segments, or for performance differences across clinically relevant conditions. For example, averaging over all frames may obscure poor performance during critical moments, and reporting unstratified metrics may hide failures in challenging scenarios such as for rare phases, low-quality recordings, presence of artifacts, or small anatomical structures, potentially diluting errors during rare but clinically significant events. Aggregating over meaningful temporal units and stratifying results by clinically relevant factors, including those affected by long-tail distributions (e.g., rare events or structures), as well as subgroup characteristics such as disease severity, surgical indication, or operator experience, enables more informative and clinically useful validation. Stratification can also improve transparency and fairness by highlighting differences in performance across subgroups. However, care must be taken in how metrics are applied and aggregated in stratified settings, as some metrics may behave non-intuitively when applied to imbalanced or small subgroups. | [funke2023metrics, kostiuchik2024surgical, zohar2020accurate, czempiel2023symphony, ross2023beyond, nwoye2023cholectrack20, nwoye2025surgitrack, huaulme2024global, hung2019experts] | Extended Data Fig. 9a |
@@
| P3.3 | Lack of contextualization of performance values | Performance values are reported without sufficient context, making it difficult to assess their clinical or practical relevance. For example, results may be presented without comparison to human performance, inter-rater agreement, or a meaningful performance threshold that defines an acceptable error rate for the clinical task. Additionally, the significance of observed differences between models may not be validated, resulting in misleading interpretations. Without such contextualization, the clinical utility of a method may remain unclear. Additionally, aspects of how performance information is communicated to end users, including the design of visualizations, thresholds, and labels, can influence clinical perception and decision-making, and should be considered in the contextualization process. | [reinke2024understanding, jogan2024quality, sylolypavan2023impact, ward2021challenges] | Extended Data Fig. 9b |
@@
| P3.4 | Lack of uncertainty analyses | Performance values are reported without conveying the uncertainty associated with the results, leading to overconfidence in the model’s performance. For example, confidence intervals, standard error, or standard deviations may be missing, concealing how much variability is present in the reported metrics, including across time or clinically relevant subgroups. | [christodoulou2024confidence, ban2021aggregating, shi2022attention, hao2023act, wang2023cascade] | Fig. 5 |
@@
| P3.5 | Insufficient reporting | Reporting of results lacks sufficient detail and transparency, making it difficult to interpret, compare, and reproduce findings. Critical aspects, such as the strategy and rationale for aggregating performance metrics, details of metric computation, or the limitations of the validation data, are often underreported or unclear. In some cases, reporting may be selective rather than merely incomplete, which can lead to overestimation of performance and misleading impressions of clinical safety or utility. Additionally, established reporting guidelines tailored to specific purposes (e.g., CONSORT-AI for clinical trials, CLAIM for medical imaging) are frequently not followed, leading to incomplete or inconsistent documentation. Another common gap is the lack of documentation regarding data modifications introduced by acquisition hardware or manufacturer-specific processing pipelines (e.g., compression, interlacing, or automatic enhancement), which can affect model performance in subtle and unquantifiable ways. | [koccak2025adherence, loftus2023artificial, navarro2023systematic, marwaha2023appraising, junger2022state, funke2023metrics] | Fig. 6 |
@@
Each of the identified pitfalls was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and to associated real-world risks (e.g., regulatory delay, compromised surgical safety). Tab. SN 2.1 and Tab. SN 2.2 provide detailed descriptions of consequences and risks.
@@
| Leakage and/or model overfitting | Leakage and model overfitting result from improper use of data during model development and may lead to misleadingly high performance estimates that do not reflect how the model will perform in real-world clinical scenarios, reducing real-world applicability. Data leakage occurs when information from the data used for testing is also present during development e.g., through shared pre-processing, improper data splitting, or data re-use, violating the assumption that test data are independent and unseen. Model overfitting refers to selecting or optimizing models based on patterns that do not generalize beyond the data used during development. It often occurs when no properly untouched test set is reserved, allowing models to perform well in the corresponding test set only rather than showing true generalizability. |
@@
| Overestimation of algorithm performance | Overestimation of algorithm performance refers to a mismatch between how well a surgical AI model is perceived to perform and how it actually behaves in clinical reality. This risk may arise from flawed or incomplete validation, misaligned performance metrics, or misinterpretation of reported results. As a consequence, models may appear more reliable, generalizable, or clinically useful than they truly are. Overestimation can lead to premature deployment, overreliance by clinicians, or inadequate oversight, increasing the risk of downstream errors, inefficiencies, or patient harm. |
@@
| Poor generalization to real-world applications | Poor generalization refers to the inability of an algorithm to maintain consistent performance across diverse real-world settings.This often results from shifts in context, technology, or clinical implementation that are not adequately captured during model development and validation. While a model may perform well in controlled settings, it may struggle when applied to different surgical procedures and teams, clinical workflows and sites, or medical devices (among others). For example, a model trained on data from one hospital may not generalize to another due to variations in equipment, imaging quality, or surgeon-specific practices. |
@@
| Selection of sub-optimal algorithms | Selection of sub-optimal algorithms refers to the risk of choosing models based on misleading or incomplete validation rather than true clinical utility. If performance metrics do not accurately reflect real-world applicability, models may be selected that perform well in development settings but lead to errors, inefficiencies, or reduced quality of care in clinical practice. Conversely, more effective algorithms may be overlooked, limiting clinical benefit and slowing progress in AI-assisted surgery. |
@@
| Lack of clinical/real-world adoption of surgical AI | Lack of clinical adoption of surgical AI refers to the failure of AI models to be integrated in real-world settings. Challenges such as misleading performance metrics, poor generalization, usability issues, or a mismatch between expected and actual clinical performance can create barriers to the integration and acceptance of surgical AI in the clinical workflow. Even if a model shows strong results in experimental settings, a lack of validation in real-world conditions - such as different surgical teams, workflows, or patient populations - can limit its adoption. If AI tools do not align with clinical needs or fail to provide tangible improvements over existing methods, they may struggle to gain acceptance and practical use in surgical settings. |
@@
Appendix SUPPL. NOTE 3 Results of the systematic review per pitfall
@@
63% of tasks specified the exact datasets used (n = 29)
@@
79% of datasets were only used once
@@
Hierarchies are not handled properly: 26.7%
@@
Model uses features that are not legitimate: 2.2%
@@
Not enough info on which model is trained with which dataset: 2.2%

### v2

@@ 9. Conflicts of Interest
| P2.6 | Non-suitability of hyperparameters from other domains | Hyperparameters or default metric configurations adopted from other domains (e.g., general computer vision) may not align with the specific underlying clinical needs. For example, in object detection, a high Intersection over Union (IoU) threshold may be appropriate in general computer vision tasks, but in surgical applications, lower IoU thresholds might be more suitable when only rough localization of objects (e.g., polyps) is needed [Tran et al., 2023]. Similarly, default hyperparameters for event detection in action recognition may not account for the variability in surgical workflows. In marker-less tool tracking (e.g., [Hein et al., 2025]), for example, thresholds that define acceptable accuracy in everyday scenarios (e.g., a few millimeters) may be insufficient in surgical contexts, where sub-millimeter precision can be clinically critical. | [Tran et al., 2023, Ramesh et al., 2023, Hein et al., 2025] | Extended Data Fig. 8b |
@@
| P2.7 | Intrinsic limitations of individual metrics | The used metrics harbor pitfalls related to their individual mathematical properties. Even if metrics are well-aligned with the clinical task, they may exhibit problematic behaviors such as sensitivity to class imbalance, non-linearity, or lack of interpretability. The suitability of any given metric should thus be analyzed in light of its known limitations. For example, accuracy may produce misleading values in highly imbalanced data sets. | [Reinke et al., 2021, 2024] | |
@@
| P3.1 | Non-independence within the test set | Failure to account for non-independence within the test set can yield misleading conclusions. Non-independence can occur when multiple frames from the same surgical video are included, or when data from the same patient, procedure, or institution are used. Aggregation, statistical analysis, and reporting need to account for this lack of independence rather than assuming independent samples. For example, if frames from the same patient are used for validation, performance metric values may appear artificially high due to strong temporal correlation. Other possible consequences include biased statistical estimates, such as underestimated uncertainty or distorted means. | [Kostiuchik et al., 2024, Funke et al., 2023, Kirtac et al., 2022] | Fig. 4 |
@@
| P3.2 | Clinically uninformative aggregation | Performance metrics are often simply aggregated, without accounting for the varying importance of different surgical phases or time segments, or for performance differences across clinically relevant conditions. For example, averaging over all frames may obscure poor performance during critical moments, and reporting unstratified metrics may hide failures in challenging scenarios such as for rare phases, low-quality recordings, presence of artifacts, or small anatomical structures, potentially diluting errors during rare but clinically significant events. Aggregating over meaningful temporal units and stratifying results by clinically relevant factors, including those affected by long-tail distributions (e.g., rare events or structures), as well as subgroup characteristics such as disease severity, surgical indication, or operator experience, enables more informative and clinically useful validation. Stratification can also improve transparency and fairness by highlighting differences in performance across subgroups. However, care must be taken in how metrics are applied and aggregated in stratified settings, as some metrics may behave non-intuitively when applied to imbalanced or small subgroups. | [Funke et al., 2023, Kostiuchik et al., 2024, Zohar et al., 2020, Czempiel, 2023, Roß et al., 2023, Nwoye et al., 2023b, Nwoye and Padoy, 2025, Huaulmé et al., 2024, Hung et al., 2019] | Extended Data Fig. 9a |
@@
| P3.3 | Lack of contextualization of performance values | Performance values are reported without sufficient context, making it difficult to assess their clinical or practical relevance. For example, results may be presented without comparison to human performance, inter-rater agreement, or a meaningful performance threshold that defines an acceptable error rate for the clinical task. Additionally, the significance of observed differences between models may not be validated, resulting in misleading interpretations. Without such contextualization, the clinical utility of a method may remain unclear. Additionally, aspects of how performance information is communicated to end users, including the design of visualizations, thresholds, and labels, can influence clinical perception and decision-making, and should be considered in the contextualization process. | [Reinke et al., 2024, Jogan et al., 2024, Sylolypavan et al., 2023, Ward et al., 2021] | Extended Data Fig. 9b |
@@
| P3.4 | Lack of uncertainty analyses | Performance values are reported without conveying the uncertainty associated with the results, leading to overconfidence in the model’s performance. For example, confidence intervals, standard errors, or standard deviations may be missing, concealing how much variability is present in the reported metrics, including across time or clinically relevant subgroups. | [Christodoulou et al., 2024, Ban et al., 2021, Shi et al., 2022, Hao et al., 2023, Wang et al., 2023] | Fig. 5 |
@@
| P3.5 | Insufficient reporting | Reporting of results lacks sufficient detail and transparency, making it difficult to interpret, compare, and reproduce findings. Critical aspects, such as the strategy and rationale for aggregating performance metrics, details of metric computation, or the limitations of the validation data, are often underreported or unclear. In some cases, reporting may be selective rather than merely incomplete, which can lead to overestimation of performance and misleading impressions of clinical safety or utility. Additionally, established reporting guidelines tailored to specific purposes (e.g., CONSORT-AI for clinical trials, CLAIM for medical imaging) are frequently not followed, leading to incomplete or inconsistent documentation. Another common gap is the lack of documentation regarding data modifications introduced by acquisition hardware or manufacturer-specific processing pipelines (e.g., compression, interlacing, or automatic enhancement), which can affect model performance in subtle and unquantifiable ways. | [Koçak et al., 2025, Loftus et al., 2023, Navarro et al., 2023, Marwaha et al., 2023, Junger et al., 2022, Funke et al., 2023] | Fig. 6 |
@@
Each of the identified pitfalls was mapped to specific potential consequences, such as introduction of biases, unreliable performance assessment, or undetected failure modes, and as to associated real-world risks (e.g., regulatory delay, compromised surgical safety). Tab. SN 2.1 and Tab. SN 2.2 provide detailed descriptions of consequences and risks.
@@
| Leakage and/or model overfitting | Leakage and model overfitting result from improper use of data during model development and may lead to misleadingly high performance estimates that do not reflect how the model will perform in real-world clinical scenarios, reducing real-world applicability. Data leakage occurs when information from the data used for testing is also present during development e.g., through shared pre-processing, improper data splitting, or data re-use, violating the assumption that test data are independent and unseen. Model overfitting refers to selecting or optimizing models based on patterns that do not generalize beyond the data used during development. It often occurs when no properly untouched test set is reserved, allowing models to perform well only in the corresponding test set rather than showing true generalizability. |
@@
| Overestimation of algorithm performance | Overestimation of algorithm performance refers to a mismatch between how well a surgical AI model is perceived to perform and how it actually behaves in clinical reality. This risk may arise from flawed or incomplete validation, misaligned performance metrics, or misinterpretation of reported results. As a consequence, models may appear more reliable, generalizable, or clinically useful than these truly are. Overestimation can lead to premature deployment, overreliance by clinicians, or inadequate oversight, increasing the risk of downstream errors, inefficiencies, or patient harm. |
@@
| Poor generalization to real-world applications | Poor generalization refers to the inability of an algorithm to maintain consistent performance across diverse real-world settings. This often results from shifts in context, technology, or clinical implementation that are not adequately captured during model development and validation. While a model may perform well in controlled settings, it may struggle when applied to different surgical procedures and teams, clinical workflows and sites, or medical devices (among others). For example, a model trained on data from one hospital may not ne generalized to another due to variations in equipment, imaging quality, or surgeon-specific practices. |
@@
| Selection of sub-optimal algorithms | Selection of sub-optimal algorithms refers to the risk of choosing models based on misleading or incomplete validation rather than true clinical utility. In case performance metrics do not accurately reflect real-world applicability, models may be selected that perform well in development settings but lead to errors, inefficiencies, or reduced quality of care in clinical practice. Conversely, more effective algorithms may be overlooked, limiting clinical benefit and slowing progress in AI-assisted surgery. |
@@
| Lack of clinical/real-world adoption of surgical AI | Lack of clinical adoption of surgical AI refers to the failure of AI models to be integrated in real-world settings. Challenges such as misleading performance metrics, poor generalization, usability issues, or a mismatch between expected and actual clinical performance can create barriers to the integration and acceptance of surgical AI in the clinical workflow. Even if a model shows strong results in experimental settings, a lack of validation in real-world conditions - such as different surgical teams, workflows, or patient populations - can limit its adoption. If AI tools do not align with clinical needs or fail to provide tangible improvements over existing methods, these may struggle to gain acceptance and practical use in surgical settings. |
@@
Appendix SUPPL. NOTE 3 Practical best practices
Appendix SUPPL. NOTE 4 Results of the systematic review per pitfall
@@
63% of tasks specified the exact datasets used
@@
79% of datasets were used only once
@@
Hierarchies were not handled properly: 26.7%
@@
Model used features that were not legitimate: 2.2%
@@
Not enough info on which model was trained with which dataset: 2.2%
@@
Appendix SUPPL. NOTE 5 Comparison of validation practices across MICCAI 2023 and 2025
To assess whether the widespread validation flaws identified in our primary literature review extend to more recent work, we performed an additional targeted screening of MICCAI 2025 surgical data science papers. This complementary analysis focused on the key pitfalls reported in the main manuscript (Sec. ’Validation flaws are widespread in common practice’) and was designed to provide an indicative comparison across conference years rather than a full replication of the original screening (as in SUPPL. NOTE 4). In this section, we present a side-by-side comparison of validation and reporting practices between MICCAI 2023 and 2025.
| Question | MICCAI 2023 | MICCAI 2025 | |
| Number of articles | 46 | 89 | +43 |
| Did the data set comprise videos (rather than static frames only)? | 74% of papers used surgical video data. | 87% of papers used surgical video data. | +13% |
| Public datasets (if any) | 79% of datasets were only used once, with 38 distinct [public] datasets identified across papers. | 53% of datasets were only used once, with 86 distinct [public] datasets identified across papers. | -26%; +48 |
| Training set size (number of videos) | The median dataset contained 37 training videos (minimum: 6 training videos). | The median dataset contained 35 training videos (minimum: 1 training videos). | -2; -5 |
| Validation set size (number of videos) | The median dataset contained 10 validation videos (minimum: 2 validation videos). | The median dataset contained 7 validation videos (minimum: 2 validation videos). | -3; 0 |
| Test set size (number of videos) | The median dataset contained 22 test videos (minimum: 2 test videos). | The median dataset contained 10 test videos (minimum: 1 test videos). | -12; -1 |
| Were data set sizes reported? | 59% of papers did not (fully) report data set sizes. | 66% of papers did not (fully) report data set sizes. | +7% |
| Was the used test set untouched? | Only 47% explicitly reported an untouched test set, while this was unclear in 37% of papers. | Only 22% explicitly reported an untouched test set, while this was unclear in 74% of papers. | -25%; +37% |
| Were metric choices justified? | Only 30% properly justified their metric choice. 20% of those justified by popularity alone. | Only 38% properly justified their metric choice. 24% of those justified by popularity alone. | +8%; +4% |
| Were clinical needs reflected in the metric choice? | In 80% of papers, it was unclear whether clinical relevance had been considered when selecting metrics. | In 75% of papers, it was unclear whether clinical relevance had been considered when selecting metrics. | -5% |
| Were algorithm properties specific to temporal data assessed? | A total of 77% of papers did not assess properties specific to temporal data. | A total of 72% of papers did not assess properties specific to temporal data. | -5% |
| Were hierarchies handled in a specific way? | Among studies involving hierarchical structures, (e.g., patient-level), only 5% explicitly accounted for their dependencies. | Among studies involving hierarchical structures, (e.g., patient-level), only 1% explicitly accounted for their dependencies. | -4% |
| Metrics | Only a single paper [2%] used a temporal consistency metric. 41 metrics were used only once [37%]. The most commonly used metric was Accuracy. | Only seven papers [8%] used a temporal metric. 93 metrics were used only once [72%]. The most commonly used metric was IoU. | +6%; +35; Accuracy IoU |
| Was the aggregation procedure described in detail? | The aggregation procedure was unclear or not described at all in 66% of papers. | The aggregation procedure was unclear or not described at all in 58% of papers. | -8% |
| Were performance values put into context? | 80% did not contextualize performance values, for example against human baseline or clinical thresholds. | 73% did not contextualize performance values, for example against human baseline or clinical thresholds. | -7% |
| Were confidence or prediction intervals reported? | Only one paper [2%] reported confidence intervals, and one [2%] reported prediction intervals. | Only one paper [1%] reported confidence intervals, and no paper [0%] reported prediction intervals. | -1%; -2% |
| Was inter-rater variability reported? | 98% of papers did not report inter-rater variability. | 95% of papers did not report inter-rater variability. | -3% |
| Did authors report adhering to common reporting guidelines? | Only a single paper [2%] reported sufficient detail to enable reproducibility. | Only three papers [3%] reported sufficient detail to enable reproducibility. | +1% |
| Were ethical aspects reported? | 78% [of papers] lacked ethical reporting. | 84% [of papers] lacked ethical reporting. | +6% |
| Were fairness-/ bias-/ equity-related aspects reported? | 89% [of papers] ignored fairness or biases. | 94% [of papers] ignored fairness or biases. | +5% |
| Were societal-/ legal-/ governance-related aspects reported? | 91% [of papers] omitted social, legal, or governance considerations. | 89% [of papers] omitted social, legal, or governance considerations. | -2% |